# Astrophysical Implications of Eccentricity in Gravitational Waves from Neutron Star-Black Hole Binaries

### v1

@@ III.2 Astrophysical implications of detecting eccentric NSBHs
In Figure 4, we plot the probability distribution over $\beta_{t}$ assuming $N_{\mathrm{ecc}}=1$ for $N_{\mathrm{obs}}=3$ and $N_{\mathrm{obs}}=11$. When $N_{\mathrm{obs}}=3$, the peak and majority of the likelihood distribution is in the unphysical region above $\beta_{t}=1$, implying a slight overabundance of eccentric mergers given the expected detection rates.
IV Discussion & Conclusion
@@
In this work we investigate the measurability and detectability of eccentricity in GW signals from merging NSBHs. We find that, while NSBHs similar to those that have been detected so far have similar minimum measurable eccentricities to binary BHs of $\mathcal{O}(0.01)$, those that exhibit significant higher-mode content may have measurable eccentricities an order of magnitude lower. We also find that the presence of non-zero precessing spin may decrease the minimum measurable eccentricity (i.e., improving our ability to measure eccentricity) for more unequal-mass NSBH mergers. By simulating a quasi-circular templated search for a realistic population of NSBH mergers produced by field triples, we find that many signals from these binaries would not be found in current detectors and pipelines, with $73\%$ having an SNR computed against a quasi-circular, aligned-spin template that is below our assumed detection threshold.
@@
This low recovery fraction has two implications: firstly, that a large population of triple-origin NSBHs could be merging in the Universe without being found by our searches; and secondly, that a single detection of an eccentric NSBH implies a high rate of NSBH mergers from triples. Indeed, we find that if we have observed only three NSBH mergers and one of them is measurably eccentric, then the likelihood for the branching fraction $\beta_{t}$ of NSBHs from triples peaks above the physical upper prior limit of $\beta_{t}=1$. On the other hand, if all 11 marginal and ambiguous NSBH candidates in the GW transient catalogue so far are counted as detections, then one measurably eccentric merger yields a $\beta_{t}$ peak below $1$.
@@
If triples are efficiently producing NSBH mergers, then they may also efficiently produce binary BH mergers. We speculate that the fraction of detectably eccentric mergers should be smaller for a population of binary BHs evolved with the same assumptions as StegmannKlencki2025, because the fractional mass loss during compact-object formation and compact-object natal kicks are smaller for BHs than NSs, so more distant tertiaries can remain bound and lead to less aggressive orbital driving. This is consistent with predictions of eccentricity distributions in binary BHs from triples [e.g., Antonini2016, Antonini2017, Liu2018, Dorozsmai2025]; however, we note that the results of these triple-evolving binary BH simulations cannot easily be compared to the results of StegmannKlencki2025, since they employ different assumptions and cover different regions of parameter space. The relative rates of binary BH and NSBH mergers from triples are also uncertain, given the lack of self-consistent simulations of triples producing both types of binary merger. For binary NSs, HamersThompson:2019:triples find that merger rates can be an order of magnitude higher than for NSBHs if natal kicks are low, but can be lower if natal kicks are high. A full population synthesis study with consistent assumptions, aiming to quantify the relationship between the NSBH merger rate, the binary BH merger rate, and the binary NS merger rate from triples, is left for future work.
@@
In the simulations of StegmannKlencki2025, for every triple-origin NSBH produced with $e_{22,10}<0.05$, there are $\approx 4$ produced with higher eccentricity, which are less likely to be detected. As shown in the right-hand panel of Figure 1, our mock search suggests that only $4\%$ of NSBHs with an eccentricity of $e_{22,10}\approx 0.3$—similar to that claimed for GW200105—would be detected with a quasi-circular aligned-spin search. Consequently, for every eccentric NSBH detected with a quasi-circular aligned-spin search, there could be many more signals with similar or higher eccentricities that are missed. This motivates the development of search strategies for inspiral-dominated signals with a wide range of total masses, eccentricities, and spin orientations.
@@
Acknowledgements
We thank Nick Loutrel, Antoni Ramos-Buades, P.J. Nee, Nihar Gupte, and Aldo Gamboa for helpful discussions. We also thank Tom Dent, Ben Patterson, Alejandra Gonzalez, and Arif Shaikh for comments on the draft. I.M.R-S acknowledges support from the Science and Technology Facilities Council Ernest Rutherford Fellowship grant number UKRI2423. M.Z. gratefully acknowledges funding from the Brinson Foundation in support of astrophysics research at the Adler Planetarium. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.

### v2

@@ 3.2 Astrophysical implications of detecting eccentric NSBHs
In Figure 4, we plot the probability distribution over $\beta_{t}$ assuming $N_{\mathrm{ecc}}=1$ for $N_{\mathrm{obs}}=3$ and $N_{\mathrm{obs}}=11$. In both cases, the peak and majority of the likelihood distribution is in the unphysical region above $\beta_{t}=1$, implying an overabundance of eccentric mergers given the expected detection rates.
4 Discussion & Conclusion
@@
In this work we investigate the measurability and detectability of eccentricity in GW signals from merging NSBHs. We find that, while NSBHs similar to those that have been detected so far have similar minimum measurable eccentricities to binary BHs of $\mathcal{O}(0.01)$, those that exhibit significant higher-mode content may have measurable eccentricities an order of magnitude lower. We also find that the presence of non-zero precessing spin may decrease the minimum measurable eccentricity (i.e., improving our ability to measure eccentricity) for more unequal-mass NSBH mergers. By simulating a quasi-circular templated search for a realistic population of NSBH mergers produced by field triples, we find that many signals from these binaries would not be found in current detectors and pipelines, with $96.1\%$ having an SNR computed against a quasi-circular, aligned-spin template that is below our assumed detection threshold.
@@
This low recovery fraction has two implications: firstly, that a large population of triple-origin NSBHs could be merging in the Universe without being found by our searches; and secondly, that a single detection of an eccentric NSBH implies a high rate of NSBH mergers from triples. Indeed, we find that if we have observed only three NSBH mergers and one of them is measurably eccentric, then the likelihood for the branching fraction $\beta_{t}$ of NSBHs from triples peaks above the physical upper prior limit of $\beta_{t}=1$. Even if all 11 marginal and ambiguous NSBH candidates in the GW transient catalogue so far are counted as detections, then one measurably eccentric merger yields a $\beta_{t}$ that is still above $1$.
@@
If triples are efficiently producing NSBH mergers, then they may also efficiently produce binary BH mergers. We speculate that the fraction of detectably eccentric mergers should be smaller for a population of binary BHs evolved with the same assumptions as Stegmann and Klencki (2025), because the fractional mass loss during compact-object formation and compact-object natal kicks are smaller for BHs than NSs, so more distant tertiaries can remain bound and lead to less aggressive orbital driving. This is consistent with predictions of eccentricity distributions in binary BHs from triples (e.g., Antonini et al., 2016, 2017; Liu and Lai, 2018; Dorozsmai et al., 2025); however, we note that the results of these triple-evolving binary BH simulations cannot easily be compared to the results of Stegmann and Klencki (2025), since they employ different assumptions and cover different regions of parameter space. The relative rates of binary BH and NSBH mergers from triples are also uncertain, given the lack of self-consistent simulations of triples producing both types of binary merger. For binary NSs, Hamers and Thompson (2019) find that merger rates can be an order of magnitude higher than for NSBHs if natal kicks are low, but can be lower if natal kicks are high. A full population synthesis study with consistent assumptions, aiming to quantify the relationship between the NSBH merger rate, the binary BH merger rate, and the binary NS merger rate from triples, is left for future work.
@@
In the simulations of Stegmann and Klencki (2025), for every triple-origin NSBH produced with $e_{22,10}<0.05$, there are $\approx 4$ produced with higher eccentricity, which are less likely to be detected. As shown in the right-hand panel of Figure 1, our mock search suggests that $1\%$ of NSBHs with an eccentricity of $e_{22,10}\approx 0.3$—similar to that claimed for GW200105—would be detected with a quasi-circular aligned-spin search. Consequently, for every eccentric NSBH detected with a quasi-circular aligned-spin search, there could be many more signals with similar or higher eccentricities that are missed. This motivates the development of search strategies for inspiral-dominated signals with a wide range of total masses, eccentricities, and spin orientations.
Acknowledgements
We sincerely thank Alex Nitz for thoroughly assessing our work, and in particular for spotting an error in our computations that led to artificially high recovered detection fractions from our searches in v1 of this manuscript. We also thank Nick Loutrel, Antoni Ramos-Buades, P.J. Nee, Nihar Gupte, and Aldo Gamboa for helpful discussions, and thank Tom Dent, Ben Patterson, Alejandra Gonzalez, and Arif Shaikh for comments on the draft. I.M.R-S acknowledges support from the Science and Technology Facilities Council Ernest Rutherford Fellowship grant number UKRI2423. M.Z. gratefully acknowledges funding from the Brinson Foundation in support of astrophysics research at the Adler Planetarium. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.